Suppose you are given the following information: Corporate Tax rate: 30% Team revenue: $20 million Team expenditure on players: $8 million Total team expenditure: $15 million You work for this franchise and are asked to estimate the taxes that will be due this year(assume the only tax is the corporate tax on revenue) . How much is due in taxes if there is no player depreciation? How much is due if we assume players depreciate 25% this year?
Answer #1

Solution:

When there is no player depreciation, Net team revenue = Total team revenue - total team expenditure

= 20 - 15 = $5 million

Corporate tax rate = 30%

Hence taxes due this year = 5*30% = $1.5 million

When there is player depreciation of 25%, the team expenditure on players will increase by 25%.

Team expenditure on players = 8*125% = $10 million.

Total team expenditure = 15 + 2 = $17 million because there is increase of expenditure of $2 million.

Net revenue = 20 - 17 = $3million

Taxes due = 3*30% = $0.9 million